# Robust Bayesian Inference for Berkson and Classical Measurement Error Models

### v2

@@ Abstract
Measurement error occurs when a covariate influencing a response variable is corrupted by noise. This can lead to misleading inference outcomes, particularly in problems where accurately estimating the relationship between covariates and response variables is crucial, such as causal effect estimation. Existing methods for dealing with measurement error often rely on strong assumptions such as knowledge of the error distribution or its variance and availability of replicated measurements of the covariates. We propose a Bayesian Nonparametric Learning framework that is robust to mismeasured covariates, does not require the preceding assumptions, and can incorporate prior beliefs about the error distribution. This approach gives rise to a general framework that is suitable for both Classical and Berkson error models via the appropriate specification of the prior centering measure of a Dirichlet Process (DP). Moreover, it offers flexibility in the choice of loss function depending on the type of regression model. We provide bounds on the generalization error based on the Maximum Mean Discrepancy (MMD) loss which allows for generalization to non-Gaussian distributed errors and nonlinear covariate-response relationships. We showcase the effectiveness of the proposed framework versus prior art in real-world problems containing either Berkson or Classical measurement errors.
@@ 1 Introduction
An often encountered problem in statistical inference is when one or more covariate variables are subject to measurement error (ME). This is known as the ME, errors-in-variables, or input uncertainty problem and often arises in medical, economic, and natural sciences when it is hard or impossible to measure quantities in the real world exactly. If ignored, ME in the covariates can result in biased parameter estimates (see Carroll et al.,, 1995). If the goal is solely to make predictions, ME may not be a significant issue as long as observations continue to be realizations of the same random variable (RV). However, in many cases, we are interested in estimating the true parameter that explains the relationship between two RVs. For example, suppose we want to estimate a parameter that represents the condition of patients based on the prevalence of a specific protein in their blood. This information can help clinicians determine the appropriate dosage of a drug that adjusts the protein levels. However, the available blood test may only provide an approximate measurement of the actual protein percentages. If this ME is ignored, the parameter estimates might give a misleading representation of the effect of the protein levels on the condition, leading to administering incorrect drug doses which could potentially be harmful to patients.
Moreover, in causal inference, a biased estimate can lead to false conclusions about the causal relationship between a covariate and the response variable. This problem has been explored in the causal inference literature by reframing it as a causal effect estimation problem with unobserved confounders (see Kuroki and Pearl,, 2014; Adams et al.,, 2019; Finkelstein et al.,, 2021). Such causal effect estimation problems are met in health and epidemiology sciences where we are interested in exposure-response relations (see Brakenhoff et al.,, 2018). Exposures, often measured with error, can include both self-reported quantities, such as patient lifestyle changes, and laboratory measurements such as serum cholesterol or the levels of aluminum in the brain (Campbell,, 2002).
@@
The ME problem has been extensively researched in both frequentist and Bayesian contexts. However, many proposed methods rely on knowing the ME distribution or variance, strong distributional assumptions, or availability of replicate measurements and instrumental variables. Our primary aim is to develop a broadly applicable Bayesian approach, robust to ME on covariates when the model doesn’t explicitly address ME and replicate measurements are unavailable. The focus is to provide informative, well-calibrated posterior densities for parameters of interest while incorporating any available expert prior beliefs on ME, without assuming that these are well-specified. If strong prior beliefs are absent, our method can still be employed by using non-informative priors or reducing the weight on priors. We introduce a Bayesian Nonparametric Learning (NPL) framework suitable for both Classical and Berkson error settings to achieve this.
1.1 Problem setting & Related work
@@
Consider two topological spaces $\mathcal{X}$ and $\mathcal{Y}$ equipped with Borel $\sigma$-algebras $\mathcal{\sigma}_{\mathcal{X}}$ and $\mathcal{\sigma}_{\mathcal{Y}}$ respectively and let $\mathcal{P}_{\mathcal{X}}$ and $\mathcal{P}_{\mathcal{Y}}$ be the space of Borel distributions on $\mathcal{X}$ and $\mathcal{Y}$ respectively. Let also $\mathcal{Z}:=\mathcal{X}\times\mathcal{Y}$ equipped with $\sigma$-algebra $\mathcal{\sigma}_{\mathcal{Z}}$ and $\mathcal{P}_{\mathcal{Z}}$ the space of Borel distributions on $\mathcal{Z}$. Assume that we have observed data $\{(w_{i},y_{i})\}_{i=1}^{n}\in\mathcal{Z}$ and wish to perform parametric inference for a model family $\{\operatorname{\mathbb{P}}_{g(\theta,x)}:x\in\mathcal{X},\theta\in\Theta\}\subset\mathcal{P}_{\mathcal{Y}}$ with regression function $g:\Theta\times\mathcal{X}\rightarrow\mathbb{R}$ and parameter space $\Theta$ such that $x\rightarrow\operatorname{\mathbb{P}}_{g(\theta,x)}(A)$ is $\sigma_{\mathcal{X}}$-measurable for all $A\in\sigma_{\mathcal{Y}}$ and all $\theta\in\Theta$. Further assume that we have RVs $W,X,N\in\mathcal{X}$ such that observations $\{w_{i}\}_{i=1}^{n}$ from $W$ are equal to the unobserved true values $\{x_{i}\}_{i=1}^{n}$ from $X$ corrupted with additive error $\{\nu_{i}\}_{i=1}^{n}$ from $N$. Assume that there exists unknown $\theta_{0}\in\Theta$ such that:
@@
for unknown $\operatorname{\mathbb{F}}_{E}^{0}$ with $\mathbb{E}_{\operatorname{\mathbb{F}}_{E}}[E]=0$. We further assume that both $N$ and $E$ are homoscedastic, i.e. they have constant variance across observations. In this work, we consider the two most common types of ME, namely Berkson (Berkson,, 1950) and Classical. The main difference lies in the assumptions of dependence between the true and noisy covariates with the measurement error. Mathematically we define:
@@
Berkson ME occurs when the user assigns the observed covariates and are hence independent of the ME. An example is when a group average ($W$) of the quantity of interest is assigned to each individual of the group whereas the true individual levels ($X$) differ by an additive error. Another example occurs when a health practitioner prescribes a fixed dose of a drug ($W$) whereas the actual quantity of the drug substance absorbed by each body ($X$) differs between individuals with the same prescription. In contrast, classical ME often occurs when the ME is a source of inaccurate recording methods such as a faulty instrument. As a result, observed measurements ($W$) randomly fluctuate around the true value ($X$) and is hence correlated with the ME. The difference between Berkson and Classical ME can also be understood from a graphical representation as illustrated in Figs. 1(a) and 1(b). In the Berkson case, $N$ directly affects $X$ whereas in the Classical case dependence flows from $N$ to $W$.
@@
Figure 1: Graphical representation of Berkson and Classical ME. Directed arrows indicate conditional dependency. The function $g(\theta,\cdot)$ indicates the regression function explaining the relationship of $Y$ conditioned on $X$. The difference between Berkson and Classical models relies on the dependency relationships between $N,W$ and $X$.
In both cases, we assume that $N$ is random and $\operatorname{\mathbb{F}}_{N}^{0}$ might not be known exactly. The goal of this work is to perform parametric inference on $\theta\in\Theta$, remaining robust with respect to ME and potential misspecifications of apriori beliefs about its distribution. Below, we outline key literature on ME, emphasizing their relevance to our work.
Numerical optimization Approaches
In the numerical optimization literature, the TLS method (Golub and Loan,, 1979; Golub and Van Loan,, 1980) has been widely used to approach ME in the covariate. This is a generalization of Ordinary Least Squares (OLS) wherein the residuals of both the response and covariate variables are minimized. The ordinary TLS algorithm has been extended to, among others, the regularised TLS (Guo and Renaut,, 2002) to induce regularisation on the set of admissible solutions and convex TLS (Malioutov and Slavov,, 2014) to deal with varying ME, known error structure and outliers. Diao et al., (2019) later improved the computational time of solving the TLS problem through fast sketching methods. This group of methods gives very useful computational tools for approaching the ME problem, however, their use in nonlinear regression problems can be challenging while they also make the underlying assumption that the errors are Normally distributed. Moreover, they fail to incorporate any prior beliefs and quantify any uncertainty related to inference outcomes. For this reason, we approach TLS from a probabilistic point of view: we incorporate it as a loss function in the Bayesian NPL framework (Lyddon et al.,, 2018), based on randomized objective functions.
$\forall i=1,\dots,n:\quad y_{i}=\theta^{T}x_{i}+\epsilon_{i},\quad\quad x_{i}=w_{i}+\nu_{i}$ (2)
and we wish to correctly estimate $\theta\in\mathbb{R}^{d_{x}\times 1}$. It is straightforward to see that if we naively use OLS to fit $\{(w_{i},y_{i})\}_{i=1}^{n}$ with the linear model, ignoring ME, the resulting estimator for $\theta$ will be biased because each $w_{i}$ is correlated with $\nu_{i}$. This bias is not present in the TLS case. Denote by $\|\cdot\|_{F}$ the Frobenius norm, then TLS corresponds to the optimization problem:
or equivalently:
The solution to the problem in (3) has been derived for real RVs using the Singular Value Decomposition (SVD) by Golub and Loan, (1979), and Golub and Van Loan, (1980) and later generalized by Van Huffel and Vandewalle, (1988) who relaxed some of the necessary numerical assumptions for the existence of a solution. A detailed solution to the problem is given in Appendix A.2. From a statistical point of view, the TLS objective has an underlying assumption of Gaussian distributed errors with (3) corresponding to maximizing the joint log-likelihood in the Classical ME case under the assumption that $\operatorname{\mathbb{F}}_{N}^{0}$ and $\operatorname{\mathbb{F}}_{E}^{0}$ are both zero mean Gaussian distributions (see Appendix A.1). Some finite sample statistical properties of TLS estimators were given by Xu et al., (2012). For a more thorough review of TLS methods, we refer the reader to Van Huffel, (2004) and Markovsky and Van Huffel, (2007).
Frequentist Methods
@@
In the statistical literature, most existing methods require knowledge of the distribution of the ME, or its variance (Cook and Stefanski,, 1994). Specifically, in Deming regression (Deming,, 1943) for Classical ME the ratio of the variances of the errors associated with the response and covariate variables is assumed to be known. For Berkson ME, Huwang and Huang, (2000) suggested an iterative re-weighted least squares for polynomial regression, assuming Gaussian distributed Berkson error. Wang, (2003) considered the more general case of nonlinear regression with Gaussian distributed errors and later generalized it to non-Gaussian error models (Wang,, 2004) through a minimum distance estimation approach.
@@
Cook and Stefanski, (1994) proposed the simulation-extrapolation method (SIMEX) which assumes knowledge of the ME variance or the ability to estimate it. Under this information, SIMEX simulates additional ME to estimate its effect on the relationship between covariate and response variables. Although SIMEX has been studied primarily for the Classical error model, some works have extended it to the case of Berkson error models (Althubaiti and Donev,, 2016) and of mixtures of Classical and Berkson errors (Carroll et al.,, 2006; Misumi et al.,, 2018).
Since the ME variance is rarely known in practice, alternative methods assume instead that we have access to repeated measurements of the mismeasured RV. These methods, usually dealing with Classical type error, (e.g. see Liu and Liang,, 1992; Li and Vuong,, 1998; Li,, 2002) treat the true values of the covariates as latent variables and use the replicate measurements to estimate them. Kato et al., (2021) showed that such techniques achieve robustness against identification failure and confidence bands for the marginal distribution of the true covariate in the Classical ME case. A widely studied method, used in both Classical and Berkson error models, is Regression Calibration (Carroll et al.,, 2006). It is based on a two-step approach, starting with the regression of $X$ on $W$ and then the substitution of the fitted model into the regression of $Y$ on $X$. Although it constitutes a very powerful method, in practice, the first step requires the existence of replicate measurements of the unobserved covariate.
Instrumental variable (IV) methods, for example Newhouse and McClellan, (1998); Bowden and Turkington, (1990) in the Classical setting, relax the above assumption but require access to observations from the so-called IV, which is correlated with the covariate, but not necessarily a replicate of it. The required conditional independence between the IV and the response variable, given the unobserved covariate, overcomes the bias induced by the correlation between the regression error and the covariate in the standard setting. The IV setting has also been generalized to a kernel-based nonparametric estimation of causal effect with Classical ME in the cause, in the presence of unobserved continuous confounding by Zhu et al., (2022). In the Berkson error case, Schennach, (2013) deals with nonparametric regression relying on the availability of an IV. Even in the cases where a candidate RV is available, it is hard to ensure that the necessary IV conditions hold.
In the nonparametric literature, the problem of Classical ME has been tackled through the deconvolution kernel methods (Fan and Truong,, 1993; Wang and Wang,, 2011) for kernel estimation of nonparametric regression functions. One of the main challenges in this method is the choice of the deconvolving kernel as these do not generally admit a closed form and can be sensitive to the selection of bandwidth. Lütkenöner, (2015) provides a family of kernels which gives rise to an analytical functional representation of the associated deconvolving kernels when the ME is normally distributed. Later, Delaigle and Hall, (2016) relaxed certain assumptions and provided a deconvolution approach for unknown, symmetric error distribution. In the same setting, Kato and Sasaki, (2018) constructed uniform confidence bands under a multiplier-bootstrap method. While the error distribution is assumed to be unknown, the method requires an auxiliary independent sample from the error distribution of the response variable. Although this group of methods has desired theoretical properties, they are usually limited to a particular function class such as sufficiently smooth functions. Nonparametric methods have also been suggested for treating Berkson-type error models. Delaigle et al., (2006) provided a flexible nonparametric approach based on trigonometric functions. The method assumes some knowledge of the ME distribution or availability of replicate measurements of the error or the covariates.
Bayesian Approaches
@@
In the Bayesian setting, most methods relax the assumptions discussed above but are often reliant on strong parametric assumptions. In the Classical error case, for example, Berry et al., (2002) restrict the distributions of the covariates, response error and ME to a Gaussian distribution and model the regression function with smoothing splines and penalized mixtures of truncated polynomial splines. Sarkar et al., (2014) proposed the use of mixtures of B-splines and mixtures induced by Dirichlet Processes (DPs) to target the more general case of heteroscedastic errors. More flexible models based on DPs and Gaussian mixtures have been also proposed by Müller and Roeder, (1997); Carroll et al., (1999). While these methods overcome the restrictions of simple parametric forms, they are usually computationally expensive to sample from as they rely on MCMC techniques. Recently, neural networks have been used in nonparametric regression with ME (Hu et al.,, 2022) to handle more generic regression functions in large-scale scenarios. In this work, the authors employ variational inference and gradient descent to overcome the computational challenges of existing MCMC methods. However, knowledge of the ME distribution type is again required. Finally, in the nonparametric Bayesian literature, Gaussian Processes (GPs) have been employed to tackle ME on the input covariate. Cervone and Pillai, (2015) adopted GP modeling for location error on the input, based on a Hybrid Monte Carlo algorithm. More recently Zhou et al., (2023) proposed a fully Bayesian approach based on a GP prior on the regression function coupled with a Dirichlet process Gaussian mixture prior on the distribution of $X$, offering theoretical guarantees on the consistency and convergence of the posterior distribution. Gaussian processes have also been employed in the Berkson error case by Zhang et al., (2017) who studied tunnel deformation detection. However, to the best of our knowledge, a more general, fully Bayesian treatment of Berkson ME has not yet been offered.
Our framework focuses on Bayesian parametric inference via a nonparametric prior on the unknown conditional distribution of $X\>|\>W$. It can be applied to any parametric model with continuous response, without assuming a well-specified ME distribution a priori. Importantly, it is suitable for both Classical and Berkson error models via a flexible specification of the nonparametric prior. In addition, from a computational point of view, the method gives rise to a parallelizable algorithm and hence overcomes computational challenges related to existing Bayesian methods.
1.2 Contributions and paper outline
We introduce a Bayesian framework for regression problems when the covariate is observed with Berkson or Classical ME. Contrary to existing methods, we do not require replicate measurements of the covariate or instrumental variables and do not assume that the specification of the ME or covariate distributions is correct. Our framework can incorporate prior beliefs about the ME distribution, such as its variance, and the degree of confidence in these. We offer a simply parallelizable algorithm, across bootstrap iterations, for sampling from the desired posterior distribution of the parameters of interest and further provide intuition on the prior specification through theoretical bounds on the generalisation error of our method.
The paper is outlined as follows. Section 2 provides the necessary background on the Bayesian NPL framework. In Section 3 we present the methodology for robust parametric inference in the ME model setting (Robust-MEM) for both Berkson and Classical error models and we suggest two loss functions based on the TLS and the MMD. In Section 4 we theoretically analyze our framework in the generic, nonlinear regression case by providing generalisation error bounds using the MMD-based loss function. Finally, in Section 5 we demonstrate strong empirical results on simulations and mental health and Economics datasets.

### v3

# Robust Bayesian Inference for Measurement Error Misspecification: The Berkson and Classical Cases

@@ Abstract
Measurement error occurs when a covariate influencing a response variable is corrupted by noise. This can lead to misleading inference outcomes, particularly in problems where accurately estimating the relationship between covariates and response variables is crucial, such as causal effect estimation. Existing methods for dealing with measurement error often rely on strong assumptions such as knowledge of the error distribution or its variance and availability of replicated measurements of the covariates. We propose a Bayesian Nonparametric Learning framework that is robust to misspecification of these assumptions and does not require replicate measurements. This approach gives rise to a general framework that is suitable for both Classical and Berkson error models via the appropriate specification of the prior centering measure of a Dirichlet Process (DP). Moreover, it offers flexibility in the choice of loss function depending on the type of regression model. We provide bounds on the generalisation error based on the Maximum Mean Discrepancy (MMD) loss which allows for generalisation to non-Gaussian distributed errors and nonlinear covariate-response relationships. We showcase the effectiveness of the proposed framework versus prior art in real-world problems containing either Berkson or Classical measurement errors.
@@ 1 Introduction
An often encountered problem in statistical inference is when one or more covariate variables are subject to measurement error (ME). This is known as the ME, errors-in-variables, or input uncertainty problem and often arises in medical, economic, and natural sciences when it is hard or impossible to measure quantities in the real world exactly. If ignored, ME in the covariates can result in biased parameter estimates (see Carroll et al.,, 1995). If the goal is solely to make predictions, ME may not be a significant issue as long as observations continue to be realizations of the same random variable (RV). However, in many cases, we are interested in estimating the true parameter that explains the relationship between two RVs. For example, suppose we want to estimate a parameter that represents patients’ condition based on the prevalence of a specific protein in their blood. This information can help clinicians determine the appropriate drug dosage that adjusts the protein levels. However, the available blood test may only provide an approximate measurement of the actual protein percentages. If this ME is ignored, the parameter estimates might give a misleading representation of the effect of the protein levels on the condition, leading to administering incorrect drug doses which could potentially be harmful to patients.
@@
The ME problem has been extensively researched in both frequentist and Bayesian contexts. However, many proposed methods rely on knowing the ME distribution or variance, strong distributional assumptions, or the availability of replicate measurements and instrumental variables. Our aim is to develop a broadly applicable Bayesian approach, robust to ME on covariates when the model doesn’t explicitly address ME and replicate measurements are unavailable. The focus is to provide informative, well-calibrated posterior densities for the parameters of interest while incorporating any available expert prior beliefs on ME, without assuming that these are well-specified. We introduce a Bayesian Nonparametric Learning (NPL) framework suitable for both Classical and Berkson error settings to achieve this.
1.1 Paper outline
The paper is outlined as follows. Section 1.3 introduces the Berkson and Classical measurement error and related work. Section 2 provides the necessary background on the Bayesian NPL framework. In Section 3 we present the methodology for robust parametric inference in the ME setting (Robust-MEM) with two loss functions based on the TLS and the MMD. In Section 4 we provide theoretical results of the generalisation error using the MMD-based loss function. Finally, in Section 5 we demonstrate strong empirical results on simulations and mental health and dietary datasets.
1.2 Notation
Consider two topological spaces $\mathcal{X}$ and $\mathcal{Y}$ equipped with Borel $\sigma$-algebras $\mathfrak{S}_{\mathcal{X}}$ and $\mathfrak{S}_{\mathcal{Y}}$ respectively and let $\mathcal{P}_{\mathcal{X}}$ and $\mathcal{P}_{\mathcal{Y}}$ be the space of Borel distributions on $\mathcal{X}$ and $\mathcal{Y}$ respectively. Let also $\mathcal{Z}:=\mathcal{X}\times\mathcal{Y}$ equipped with $\sigma$-algebra $\mathfrak{S}_{\mathcal{Z}}$ and $\mathcal{P}_{\mathcal{Z}}$ the space of Borel distributions on $\mathcal{Z}$. Throughout the paper we use the notation $\times$ for products of measures such that $\operatorname{\mathbb{P}}=\operatorname{\mathbb{P}}_{X}\times\operatorname{\mathbb{P}}_{Y\mid X}$ denotes the joint law on $\mathcal{Z}$ built from the marginal $\operatorname{\mathbb{P}}_{X}$ on $\mathcal{X}$ and the conditional law $\operatorname{\mathbb{P}}_{Y\mid X}$ on $\mathcal{Y}$, given $X$. Concretely, for measurable sets $A\in\mathfrak{S}_{\mathcal{X}}$ and $B\in\mathfrak{S}_{\mathcal{Y}}$,
$\operatorname{\mathbb{P}}(A\times B)\;=\;\int_{A}\operatorname{\mathbb{P}}_{Y\mid X}(x,B)\,\operatorname{\mathbb{P}}_{X}(dx).$
[…]
1.3 Berkson and Classical Measurement Error
@@
Assume that we have observed data $\{(w_{i},y_{i})\}_{i=1}^{n}\in\mathcal{Z}$ and wish to perform parametric inference for a model family $\mathcal{P}_{\Theta}:=\{\operatorname{\mathbb{P}}_{g(\theta,x)}:x\in\mathcal{X},\theta\in\Theta\}\subset\mathcal{P}_{\mathcal{Y}}$ with regression function $g:\Theta\times\mathcal{X}\rightarrow\mathbb{R}$ and parameter space $\Theta$ such that $x\rightarrow\operatorname{\mathbb{P}}_{g(\theta,x)}(A)$ is $\mathfrak{S}_{\mathcal{X}}$-measurable for all $A\in\mathfrak{S}_{\mathcal{Y}}$ and all $\theta\in\Theta$. Further assume that we have RVs $W,X,N\in\mathcal{X}$ such that observations $\{w_{i}\}_{i=1}^{n}$ from $W$ are equal to the unobserved true values $\{x_{i}\}_{i=1}^{n}$ from $X$ corrupted with additive error $\{\nu_{i}\}_{i=1}^{n}$ from $N$. We assume that there exists unknown $\theta_{0}\in\Theta$ such that:
@@
for unknown probability measure $\operatorname{\mathbb{F}}_{E}^{0}$ $\in\mathcal{P}_{\mathcal{Y}}$ with $\mathbb{E}_{\operatorname{\mathbb{F}}_{E}^{0}}[E]=0$. We further assume that both $N$ and $E$ are homoscedastic, i.e. they have constant variance across observations and we denote their variances as $\sigma_{\nu}^{2}:=\text{Var}_{\mathbb{F}_{N}^{0}}[N]$ and $\sigma_{\epsilon}^{2}:=\text{Var}_{\operatorname{\mathbb{F}}_{E}^{0}}[E]$ respectively. In this work, we consider the two most common types of ME, namely Berkson (Berkson,, 1950) and Classical. The main difference lies in the assumptions of dependence between the true and noisy covariates with the measurement error. Mathematically:
@@
Figure 1: Graphical representation of Berkson and Classical ME. Directed arrows indicate conditional dependency. The regression function explaining the relationship of Y|XY\>|\>X is denoted by $g(\theta,\cdot)$. The two models differ on the dependency relationships between $N,W$ and $X$.
Berkson ME occurs when the user assigns the observed covariates and they are hence independent of the ME. An example is when a group average ($W$) of the quantity of interest is assigned to each individual of the group, whereas the true individual levels ($X$) differ by an additive error. Another example occurs when a health practitioner prescribes a fixed dose of a drug ($W$) whereas the actual quantity of the drug substance absorbed by each body ($X$) differs between individuals with the same prescription. In contrast, classical ME often occurs when the ME is a source of inaccurate recording methods such as a faulty instrument. As a result, observed measurements ($W$) randomly fluctuate around the true value ($X$) and is hence correlated with the ME. For example, Classical ME often occurs in effect estimation problems met in health and epidemiology sciences. In these problems, we are interested in exposure-response relations (see Brakenhoff et al.,, 2018). Exposures, often measured with error, can include both self-reported quantities, such as patient lifestyle changes, and laboratory measurements such as serum cholesterol or the levels of aluminium in the brain (Campbell,, 2002).
The difference between Berkson and Classical ME can also be understood from a graphical representation as illustrated in Figs. 1(a) and 1(b). In the Berkson case, $N$ directly affects $X$ whereas in the Classical case dependence flows from $N$ to $W$. In both cases, we assume that $N$ is random and $\operatorname{\mathbb{F}}_{N}^{0}$ might not be precisely known. We aim to perform parametric inference on $\theta\in\Theta$, ensuring robustness to ME and potential misspecifications of a priori beliefs about its distribution. Below, we outline key literature on ME, emphasising its relevance to our work (see Appendix B for further review).
Total Least Squares for ME
The TLS method (Golub and Loan,, 1979; Golub and Van Loan,, 1980) has been widely used to approach ME in the covariate. This is a generalisation of Ordinary Least Squares (OLS) wherein the residuals of both the response and covariate variables are minimised. These methods give useful computational tools for the ME problem; however, their use in nonlinear regression problems can be challenging. From a statistical point of view, TLS assumes Gaussian-distributed errors, corresponding to maximising the joint log-likelihood in the Classical ME case under the assumption that $N$ and $E$ follow zero-mean Gaussians. More details and a solution to the TLS problem can be found in Appendix C.1, C.2.
Frequentist Methods for ME
@@
Most methods require knowledge of the distribution of the ME, or its variance. Deming regression (Deming,, 1943), introduced for Classical ME, assumes that the ratio of the regression and measurement error variances, i.e. $\sigma^{2}_{\epsilon}/\sigma^{2}_{\nu}$, is known. For Berkson ME, Huwang and Huang, (2000) suggested an iterative re-weighted least squares for polynomial regression, assuming Gaussian distributed error. Wang, (2003, 2004) considered the more general case of nonlinear regression with Gaussian and non-Gaussian distributed errors through a minimum distance estimation approach.
@@
The simulation-extrapolation method (SIMEX), suggested by Cook and Stefanski, (1994), assumes knowledge of the ME variance or the ability to estimate it. SIMEX simulates additional ME to estimate the relationship between covariate and response variables. It has been studied for both Classical and Berkson error models (Althubaiti and Donev,, 2016) as well as mixtures of Classical and Berkson errors (Carroll et al.,, 2006; Misumi et al.,, 2018), while it has also been generalised to semi-parametric models (Apanasovich et al.,, 2009). Since the ME variance is rarely known in practice, alternative methods for its estimation are often employed. For example, Kekeç and Van Keilegom, (2022) suggested using an expansion of Bernstein polynomials to identify the variance of Gaussian, additive, Classial ME.
Alternative methods for Classical ME, assume access to repeated measurements of the mismeasured RV or instrumental variables (IV) (e.g. see Liu and Liang,, 1992; Li,, 2002; Bowden and Turkington,, 1990). In the former case, true covariates are treated as latent, and replicate measurements are used to estimate them. In contrast, IV methods use the IV variable, which is often difficult to identify, to overcome the bias induced by Classical ME. In the nonparametric literature, addressing classical ME often involves deconvolution kernel methods (Fan and Truong,, 1993; Wang and Wang,, 2011) for kernel-based estimation of regression functions. Two key challenges lie in selecting the deconvolving kernel, which typically lacks a closed form and can be sensitive to bandwidth choice and the reliance on sufficiently smooth functions.
Bayesian Approaches to ME
@@
Most Bayesian methods relax the assumptions discussed above but are often reliant on strong parametric assumptions. In the Classical error case, Berry et al., (2002) restrict the distributions of the covariates, response error and ME to Gaussian distributions and model the regression function with smoothing splines and penalized mixtures of truncated polynomial splines. Sarkar et al., (2014) proposed the use of mixtures of B-splines and mixtures induced by Dirichlet Processes (DPs) to target the more general case of heteroscedastic errors. More recently, Brown, (2024) researched the geometric ergodicity of Gibbs samplers for multivariate Berkson and Classical ME, however the treatment is limited to Gaussian ME. More flexible models based on DPs and Gaussian mixtures have also been proposed by Müller and Roeder, (1997); Carroll et al., (1999). While these overcome the simplicity of parametric forms, they rely on computationally expensive MCMC techniques. Recently, neural networks have been used in nonparametric regression with ME (Hu et al.,, 2022) to handle more generic regression functions in large-scale scenarios. However, knowledge of the ME distribution type remains necessary. In the nonparametric Bayesian literature, Gaussian Processes (GPs) have been used to address ME on the input covariate. For example, Cervone and Pillai, (2015) used GP modelling for location error on the input, employing a Hybrid Monte Carlo algorithm. Recently, Zhou et al., (2023) proposed a fully Bayesian approach based on GP and Dirichlet process Gaussian mixture priors. GPs have also been applied in Berkson error scenarios by Zhang et al., (2017) in their study on tunnel deformation detection, but a general, fully Bayesian treatment of Berkson ME has not yet been offered.
Our framework presents a Bayesian approach for regression under Berkson or Classical ME, eliminating the need for covariate replicate measurements or instrumental variables. Unlike existing methods, it doesn’t assume precise ME or covariate distribution specifications. It can incorporate prior beliefs about the ME distribution, such as its variance, and the degree of confidence in these, through a nonparametric prior on the unknown conditional distribution of X|WX\>|\>W.